OpenStudy (hmayberry):

The coordinates of the vertices of a polygon are (0,3), (2,3) (2,0), (1,-4) and (-2,-1).  What is the perimeter of the polygon to the nearest tenth of a unit? 

OpenStudy (hmayberry):

@3mar

OpenStudy (3mar):

Well, I am here.

563blackghost (563blackghost):

Find the distance between each point. \(\huge\bf{d=\sqrt{(2-0)^{2}+(3-3)^{2}}}\) \(\huge\bf{d=\sqrt{(2-2)^{2}+(0-3)^{2}}}\) \(\huge\bf{d=\sqrt{(1-2)^{2}+(-4-0)^{2}}}\) \(\huge\bf{d=\sqrt{(-2-0)^{2}+(-1-3)^{2}}}\)

563blackghost (563blackghost):

Add each of these up to find the perimeter.

OpenStudy (3mar):

Two ways: 1: plot all these points and you can see the shape (square, rec, ....) 2: calculate the distance between every two points (sides' lengths) then sum their lengths which one are you familiar with?

OpenStudy (hmayberry):

both

OpenStudy (3mar):

@hmayberry 563blackghost did the job very well. So you can follow what she wrote and add them up to get the perimeter!

OpenStudy (hmayberry):

Okay thank you so much

563blackghost (563blackghost):

Thank you @3mar :)

OpenStudy (3mar):

You are welcome, sister!

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)